# Boaz Moerman

### v1

@@ 3.4. Geometric interpretation of the constant
Before giving the proof of Theorem 3.3, we give a geometric interpretation for the leading constant obtained in the theorem in the case where $L$ is adjoint rigid with respect to $(X,M)$, by interpreting the constant as an adelic integral. In particular, we show that the leading constant in Theorem 3.3 agrees with [Moe25conjecture, Conjecture 1.4].
@@
where $H_{a((X,M),L)L+K_{X}}$ is the toric height corresponding to the $\mathbb{Q}$-divisor class $a((X,M),L)L+K_{X}$ as defined in [BaTs95, Definition 2.1.7], and the measure $\tau_{(X,M^{\circ})}$ is the measure from [Moe25conjecture, §8] defined using the toric metric on the canonical divisor class $K_{X}$ as in [BaTs95, Theorem 2.1.6]. Furthermore, this agrees with [Moe25conjecture, Conjecture 1.4].
In particular, if $(X,M)$ is a pair corresponding to Campana points, then the constant is compatible with the prediction in [PSTVA21, §3.3].
@@
Strictly speaking, Conjecture [Moe25conjecture, Conjecture 1.4] only applies to proper pairs. However, every term in the conjecture straightforwardly generalizes to quasi-proper pairs, so we can still compare Theorem 3.3 with the conjecture for such pairs in general.
@@
where $\zeta$ is the Riemann zeta function. This implies that the convergence factors are simply given by $\lambda_{p}^{-1}=(1-p^{-1})^{b(\mathbb{Q},(X,M),L)}$ and $L^{*}(1,\operatorname{Pic}(X_{\overline{Q}},M^{\circ})=1$. In particular, $\tau_{(X,M^{\circ})}=\tau_{X,\infty}\times\prod_{p\,\mathrm{prime}}(1-p^{-1})^{b((X,M),L)}\tau_{X,p}$, where the measures $\tau_{X,\infty}$ and $\tau_{X,p}$ are the local measures on $X(\mathbb{R})$ and $X(\mathbb{Q}_{p})$ as in [CLTs10, §2.1.8] induced by the toric metric on the canonical divisor class $K_{X}$ as defined in [BaTs95, Theorem 2.1.6]. The $\mathbb{Q}$-divisor class $a((X,M),L)L+K_{X}$ on $X$ is represented by the $\mathbb{Q}$-divisor class $a((X,M),L)D+D_{X}$, where $D$ is as in Theorem 3.3 and $D_{X}=-\sum_{i=1}^{n}D_{i}$. For every prime number $p$ not dividing $S$, we write $\delta_{M,p}$ for the indicator function of $p$-adic $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$. For a prime number $p$ dividing $S$, we instead set $\delta_{M,p}=1$. The adelic integral is equal to the product
@@
where $H_{a((X,M),L)L+K_{X}}=\prod_{v\in\Omega_{\mathbb{Q}}}H_{a((X,M),L)D+D_{X},v}$ as in [BaTs95, Definition 2.1.5]. We will now verify that $C_{\infty}$ is equal to the real integral and that $C_{p}$ is equal to the $p$-adic integral, for each prime number $p$. We will first start with the archimedean place $\infty$.
For a maximal cone $\sigma\in\Sigma$, let $C_{\sigma}(\mathbb{R})\subset X(\mathbb{R})$ be the subset as defined in [Sal98, Notation 9.1]. As shown in the proof of [Sal98, Lemma 9.10], $C_{\sigma}(\mathbb{R})$ is the set of all $(x_{1}:\dots:x_{n})\in X(\mathbb{R})$ such that $|x_{1}|,\dots,|x_{n}|\leq 1$ and $x_{i}=1$ for all $i\in\{1,\dots,n\}$ with $\rho_{i}\not\subset\sigma$. In particular we can identify $C_{\sigma}(\mathbb{R})$ with $0,1]^{d}$. Under this identification, the measure $\tau_{X,\infty}$ corresponds to the Lebesgue measure on $0,1]^{d}$ as shown in the proof of [Sal98, Proposition 9.16]. By construction, we have $X(\mathbb{R})=\bigcup_{\sigma\in\Sigma_{\max}}C_{\sigma}(\mathbb{R})$ and the proof of [Sal98, Proposition 9.16] implies that $C_{\sigma}(\mathbb{R})\cap C_{\sigma^{\prime}}(\mathbb{R})$ for any two maximal cones $\sigma\neq\sigma^{\prime}$. Finally, for $P=(x_{1}:\dots:x_{n})\in C_{\sigma}(\mathbb{R})$, the height is simply given $H_{a((X,M),L)D+D_{X},\infty}(P)=\prod_{\begin{subarray}{c}i=1\\ \rho_{i}\subset\sigma\end{subarray}}^{n}x_{i}^{1-a_{i}}$. Thus we obtain
@@
Let $p$ be a prime. By [Sal98, Proposition 9.14], the measure $\tau_{X,p}$ restricts to the Haar measure on $\mathbb{A}^{d}(\mathbb{Z}_{p})=\mathbb{Z}_{p}^{d}$ for all toric subschemes $\mathbb{A}^{d}\subset\mathcal{X}$. In particular it follows that for all $\mathbf{m}\in\mathfrak{M}_{\textnormal{red}}$, the set
@@
The compatibility with the conjecture in [PSTVA21] follows from this identity together with the straightforward identity
@@
where $U^{\prime}$ is the complement of the support of $a((X,M),L)L+K_{X}+D_{\mathbf{m}}$, $\tau_{U^{\prime}}$ is the Tamagawa measure as in [CLTs10, Definition 2.8] and $\tau_{U^{\prime},D_{\mathbf{m}}}=\frac{\tau_{U^{\prime}}}{H_{D_{\mathbf{m}}}}$.
Finally, we verify that the leading constant agrees with the prediction in [Moe25conjecture, Conjecture 1.4]. Note that the adelic integral $\int_{x\in(\mathcal{X},\mathcal{M})(\mathbf{A}_{\mathbb{Z}[1/S]})}\frac{1}{H_{a((X,M),L)L+K_{X}}(x)}\,\mathrm{d}\tau_{(X,M^{\circ})}$ is equal to $\hat{\tau}(1)$ from [Moe25conjecture, Definition 8.3], where $1$ is the trival Brauer class. We will show that $\hat{\tau}(b)=0$ for every nontrivial Brauer class $b\in\operatorname{Br}_{1}(X,M)/\operatorname{Br}K$.
Using [Moe25conjecture, Lemma 7.8], we can identify $\operatorname{Br}_{1}(X,M^{\circ})/\operatorname{Br}\mathbb{Q}$ with the group of automorphic characters $\chi$ on $U\cong\mathbb{G}_{m}^{d}$ such that the induced character $(\chi_{\mathbf{m}})_{\mathbf{m}\in\Gamma_{M^{\circ}}}$ on $\mathbb{G}_{m}^{\Gamma_{M^{\circ}}}$ is identically $1$. For a place $v$, let $\mathcal{K}_{v}$ be the maximal compact subgroup of $\mathbb{G}_{m}(K_{v})^{d}$: i.e. for each prime number $p$ we have $\mathcal{K}_{p}=\mathbb{G}_{m}(\mathbb{Z}_{p})^{d}$ and $\mathcal{K}_{\infty}=\{\pm 1\}^{d}\subset\mathbb{G}_{m}(\mathbb{R})^{d}$. For any place $v\in\Omega_{\mathbb{Q}}$ given element $x\in(\mathcal{X},\mathcal{M})(\mathcal{O}_{v})$ and any element $t\in\mathcal{K}_{v}$, $tx$ also lies in $(\mathcal{X},\mathcal{M})(\mathcal{K}_{v})$ and
@@
which implies that $\hat{\tau}(\chi)$ is zero unless $\chi$ is identically $1$ on where $\mathcal{K}=\prod_{v\in\Omega_{\mathbb{Q}}}\mathcal{K}_{v}$. Therefore, in order to determine the Tamagawa constant, we only need to consider characters in the group $B$, where $B\subset\operatorname{Br}_{1}(X,M)/\operatorname{Br}\mathbb{Q}$ is the subgroup of characters which are trivial on $\mathcal{K}$. By class field theory [Neu99, Ch. VI, §6; Ch. VII, §6], the group of automorphic characters on $\mathbb{G}_{m}$ which are both torsion and trivial on the maximal compact subgroup is isomorphic to the class group of $\mathbb{Q}$. Thus as the class group of $\mathbb{Q}$ is trivial, $B$ only contains the trivial character, proving that only the trivial character contributes to the leading constant in [Moe25conjecture, Conjecture 1.4]. ∎
@@
[Bre01Sum, Théorème 1] Let $f\colon\mathbb{N}^{k}\rightarrow\mathbb{R}$ be a nonnegative arithmetic function and let $F$ be the associated Dirichlet series
@@
Theorem 3.10 has a typo in its original formulation [Bre01Sum, Théorème 1], as it requires the the upper bound in (P3) to be satisfied on the whole of $\mathscr{D}(\delta_{1}-\epsilon^{\prime})$. This is a stronger assumption than necessary and desired, as it would imply that $H$ is bounded on $\mathbb{R}_{>0}^{k}$, which would exclude the original application of the theorem in [Bre01Toric].
@@ 3.5.2. Choice of $\boldsymbol{\alpha}$
Recall that a linear form $N\rightarrow\mathbb{Z}$ is irreducible if it is not a positive multiple of another linear form. For $\tau$ a facet of $\sigma$, we write $u_{\sigma,\tau}\in N^{\vee}$ for the unique irreducible linear form which is zero on $\tau$ and positive on $\sigma\setminus\tau$. This linear form is simply given as $u_{\sigma,\tau}=n^{*}_{\rho_{i}}$, where $\rho_{i}$ is the unique ray in $\sigma(1)\setminus\tau(1)$. We also write $a_{\sigma,\tau}=a_{i}$. By [Sal98, Lemma 8.9(i)] there is a unique maximal cone $\sigma^{\prime}\neq\sigma$ containing $\tau$, which satisfies $\tau=\sigma^{\prime}\cap\sigma$. Since $u_{\sigma^{\prime},\tau}$ is also irreducible and zero on $\tau$, we see $u_{\sigma^{\prime},\tau}=\pm n_{\rho_{i}}^{*}$. Since there are $n_{1}\in\sigma\setminus\tau$, $n_{2}\in\sigma^{\prime}\setminus\tau$ such that $n_{1}+n_{2}\in\tau$, we see $u_{\sigma^{\prime},\tau}=-n_{\rho_{i}}^{*}=-u_{\sigma,\tau}$.
@@ 3.5.3. Choice of linear forms
Now, as in de la Bretèche’s work [Bre01Toric, §4.3], we will use the upper bound
@@
valid for all $\epsilon>0$ and $T>0$, which follows from [Ten15, Theorem II.3.8]. By shrinking $\delta_{1}$ if necessary, we can assume that $G$ extends to a holomorphic function on the topological closure of $\boldsymbol{\alpha}+\mathscr{D}(\delta_{1})$. Now since $G(\mathbf{s})$ is bounded on $\boldsymbol{\alpha}+\mathscr{D}(\delta_{1})$, this implies that condition (P3) in Theorem 3.10 is satisfied with $\delta_{2}=1/3$.
@@ 3.5.4. Determining the rank
where $A$ is the matrix such that the coefficient in position $(i,j)$ is $\langle\mu_{D^{\prime}}(\sigma_{j}),n_{\rho_{i}}\rangle$. Since a nef divisor on a toric variety is globally generated by [CLS11, Theorem 6.3.12], we see by [CLS11, Theorem 6.1.7] that $\operatorname{rank}(A)=\dim P_{L}$, where $P_{L}$ is the polyhedron associated to $L$ as defined in [CLS11, §4.3]. By [CLS11, Lemma 9.3.9] we also have $\dim P_{L}=\dim(X)$, as $L$ is big. The vector space $V$ is contained in the kernel of $\operatorname{Div}_{T}(X)_{\mathbb{Q}}\rightarrow\operatorname{Pic}(X)_{\mathbb{Q}}$. By [CLS11, Theorem 4.2.1.] this kernel has dimension $\dim(X)$, but $\operatorname{rank}(l^{(1)},\dots,l^{(n)})=\dim(X)+1$ implies $V$ has dimension $\dim(X)$ as well, so $V$ is equal to the kernel. ∎
@@
The constant $C_{\infty}$ in Theorem 3.3 does not depend on the choice of the integers $d_{1},\dots,d_{n}$ determining the pair $(X,\overline{M^{\circ}})$ as in Notation 2.27. Therefore, we will assume that the integers are chosen large enough to ensure $a(\mathbb{Q},(X,M),L)=a(\mathbb{Q},(X,\overline{M^{\circ}}),L)$. Note that such integers exist as $(X,M^{\circ})$ is quasi-proper with respect to $L$.
@@
[Bre01Sum, Théorème 2(ii), Remarques (ii)] In the setting of Theorem 3.10, assume that the following additional conditions are satisfied:
@@
Now it remains to estimate the volume of the set $D(B)$. In order to simplify notation, we assume $M=M^{\circ}$, which we can do without loss of generality as the definition of $D(B)$ only depends on $(X,M^{\circ})$ and not on $(X,M)$. The set $D(B)$ is a generalization of the set $D(B)$ defined by Salberger [Sal98, Notation 11.28] in his study of rational points on split toric varieties, and we will use the same approach he used to estimate its volume.
We regard $D(B)$ as a closed subset of the real locus of the universal torsor of $(X,M)$, where the universal torsor is as in Definition 2.30. As we assumed $M=M^{\circ}$ and $a((X,M),L)=1$, we have $\operatorname{pr}^{*}_{M}D=-D_{(X,M)}$. In order to estimate the volume of $D(B)$, Salberger splits it up as $D(B)=\cup_{\sigma\in\Sigma_{\max}}D(B,\sigma)$ using what he calls the toric canonical splitting [Sal98, Notation 11.31]. We will similarly split up $D(B)$, but we will use the splitting induced by the fan $\Sigma_{\overline{M}}$ as given in Notation 2.27, rather than $\Sigma$. The fan $\Sigma_{\overline{M}}$ has the property that for all $\mathbf{m}\in\Gamma_{M}$, the ray spanned by an element $\phi(\mathbf{m})\in N$ lies in $\Sigma_{\overline{M}}$, and all rays in $\Sigma_{\overline{M}}$ are of this form. This property will aid in computing $D(B)$.
@@
As in [Sal98, Proposition 11.22], we can describe when a point lies in $C_{M,\sigma,0}(\mathbb{R})$. The cone $\sigma$ contains exactly $d=\dim X$ rays. Let $d_{1}$ be the number of rays in $\sigma$ which lie in $\Sigma_{M}$. Let $r=\#\Sigma_{M(\sigma)}(1)-d$ be the number of raysin $\Sigma_{M(\sigma)}$ which lie outside of $\sigma$. Order the rays $\rho_{1},\dots,\rho_{r+d}$ in $\Sigma_{M(\sigma)}$ such that $\rho_{r+1},\dots,\rho_{r+d_{1}}$ are the rays in both $\sigma$ and $\Sigma_{M}$ and $\rho_{r+1},\dots,\rho_{r+d}$ are the rays in $\sigma$. Let $\mathbf{m}_{1},\dots,\mathbf{m}_{r+d}\in\Gamma_{M(\sigma)}$ be the elements corresponding to the rays $\rho_{1},\dots\rho_{r+d}$, and set $n^{(i)}=\phi(\mathbf{m}_{r+i})$ for $i=1,\dots,d$. As $n^{(1)},\dots,n^{(d)}$ are integer multiples of the ray generators of $\rho_{r+1},\dots,\rho_{r+d}$, they freely generate a finite-index sublattice $N_{\sigma}$ of $N$. Let $(\mu^{(1)},\dots,\mu^{(d)})$ be the corresponding dual $\mathbb{Z}$-basis of $N_{\sigma}^{\vee}\supset N^{\vee}$ and set
@@
The proof is identical to the proof of [Sal98, Proposition 11.22]. ∎
@@
Consequently, we find the following analogue of [Sal98, Equation (11.37)]
@@
obtained by evaluating at the anticanonical class $K_{(X,M)}$ of $(X,M)$. As the volume of the fibre above any $y\in[0,\infty)$ is equal to $\operatorname{Pic}(X,M)_{\mathrm{torsion}}\alpha_{\mathrm{Peyre}}((X,M),L)y^{b(\mathbb{Q},(X,M),L)-1}$, where $\alpha_{\mathrm{Peyre}}((X,M),L)$ is the variant of the $\alpha$-constant as given in [Moe25conjecture, Remark 8.6], integrating along the fibres of $\lambda$ gives
@@
Now the formula [Bar93, Example 2.1] for the exponential integral over a cone gives

### v2

@@ 3.4. Geometric interpretation of the constant
Before giving the proof of Theorem 3.3, we give a geometric interpretation for the leading constant obtained in the theorem in the case where $L$ is adjoint rigid with respect to $(X,M)$, by interpreting the constant as an adelic integral. In particular, we show that the leading constant in Theorem 3.3 agrees with [24, Conjecture 1.4].
@@
where $H_{a((X,M),L)L+K_{X}}$ is the toric height corresponding to the $\mathbb{Q}$-divisor class $a((X,M),L)L+K_{X}$ as defined in [5, Definition 2.1.7], and the measure $\tau_{(X,M^{\circ})}$ is the measure from [24, §8] defined using the toric metric on the canonical divisor class $K_{X}$ as in [5, Theorem 2.1.6]. Furthermore, this agrees with [24, Conjecture 1.4].
In particular, if $(X,M)$ is a pair corresponding to Campana points, then the constant is compatible with the prediction in [29, §3.3].
@@
Strictly speaking, Conjecture [24, Conjecture 1.4] only applies to proper pairs. However, every term in the conjecture straightforwardly generalizes to quasi-proper pairs, so we can still compare Theorem 3.3 with the conjecture for such pairs in general.
@@
where $\zeta$ is the Riemann zeta function. This implies that the convergence factors are simply given by $\lambda_{p}^{-1}=(1-p^{-1})^{b(\mathbb{Q},(X,M),L)}$ and $L^{*}(1,\operatorname{Pic}(X_{\overline{Q}},M^{\circ})=1$. In particular, $\tau_{(X,M^{\circ})}=\tau_{X,\infty}\times\prod_{p\,\mathrm{prime}}(1-p^{-1})^{b((X,M),L)}\tau_{X,p}$, where the measures $\tau_{X,\infty}$ and $\tau_{X,p}$ are the local measures on $X(\mathbb{R})$ and $X(\mathbb{Q}_{p})$ as in [10, §2.1.8] induced by the toric metric on the canonical divisor class $K_{X}$ as defined in [5, Theorem 2.1.6]. The $\mathbb{Q}$-divisor class $a((X,M),L)L+K_{X}$ on $X$ is represented by the $\mathbb{Q}$-divisor class $a((X,M),L)D+D_{X}$, where $D$ is as in Theorem 3.3 and $D_{X}=-\sum_{i=1}^{n}D_{i}$. For every prime number $p$ not dividing $S$, we write $\delta_{M,p}$ for the indicator function of $p$-adic $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$. For a prime number $p$ dividing $S$, we instead set $\delta_{M,p}=1$. The adelic integral is equal to the product
@@
where $H_{a((X,M),L)L+K_{X}}=\prod_{v\in\Omega_{\mathbb{Q}}}H_{a((X,M),L)D+D_{X},v}$ as in [5, Definition 2.1.5]. We will now verify that $C_{\infty}$ is equal to the real integral and that $C_{p}$ is equal to the $p$-adic integral, for each prime number $p$. We will first start with the archimedean place $\infty$.
For a maximal cone $\sigma\in\Sigma$, let $C_{\sigma}(\mathbb{R})\subset X(\mathbb{R})$ be the subset as defined in [30, Notation 9.1]. As shown in the proof of [30, Lemma 9.10], $C_{\sigma}(\mathbb{R})$ is the set of all $(x_{1}:\dots:x_{n})\in X(\mathbb{R})$ such that $|x_{1}|,\dots,|x_{n}|\leq 1$ and $x_{i}=1$ for all $i\in\{1,\dots,n\}$ with $\rho_{i}\not\subset\sigma$. In particular we can identify $C_{\sigma}(\mathbb{R})$ with $0,1]^{d}$. Under this identification, the measure $\tau_{X,\infty}$ corresponds to the Lebesgue measure on $0,1]^{d}$ as shown in the proof of [30, Proposition 9.16]. By construction, we have $X(\mathbb{R})=\bigcup_{\sigma\in\Sigma_{\max}}C_{\sigma}(\mathbb{R})$ and the proof of [30, Proposition 9.16] implies that $C_{\sigma}(\mathbb{R})\cap C_{\sigma^{\prime}}(\mathbb{R})$ for any two maximal cones $\sigma\neq\sigma^{\prime}$. Finally, for $P=(x_{1}:\dots:x_{n})\in C_{\sigma}(\mathbb{R})$, the height is simply given $H_{a((X,M),L)D+D_{X},\infty}(P)=\prod_{\begin{subarray}{c}i=1\\ \rho_{i}\subset\sigma\end{subarray}}^{n}x_{i}^{1-a_{i}}$. Thus we obtain
@@
Let $p$ be a prime. By [30, Proposition 9.14], the measure $\tau_{X,p}$ restricts to the Haar measure on $\mathbb{A}^{d}(\mathbb{Z}_{p})=\mathbb{Z}_{p}^{d}$ for all toric subschemes $\mathbb{A}^{d}\subset\mathcal{X}$. In particular it follows that for all $\mathbf{m}\in\mathfrak{M}_{\textnormal{red}}$, the set
@@
The compatibility with the conjecture in [29] follows from this identity together with the straightforward identity
@@
where $U^{\prime}$ is the complement of the support of $a((X,M),L)L+K_{X}+D_{\mathbf{m}}$, $\tau_{U^{\prime}}$ is the Tamagawa measure as in [10, Definition 2.8] and $\tau_{U^{\prime},D_{\mathbf{m}}}=\frac{\tau_{U^{\prime}}}{H_{D_{\mathbf{m}}}}$.
Finally, we verify that the leading constant agrees with the prediction in [24, Conjecture 1.4]. Note that the adelic integral $\int_{x\in(\mathcal{X},\mathcal{M})(\mathbf{A}_{\mathbb{Z}[1/S]})}\frac{1}{H_{a((X,M),L)L+K_{X}}(x)}\,\mathrm{d}\tau_{(X,M^{\circ})}$ is equal to $\hat{\tau}(1)$ from [24, Definition 8.3], where $1$ is the trival Brauer class. We will show that $\hat{\tau}(b)=0$ for every nontrivial Brauer class $b\in\operatorname{Br}_{1}(X,M)/\operatorname{Br}K$.
Using [24, Lemma 7.8], we can identify $\operatorname{Br}_{1}(X,M^{\circ})/\operatorname{Br}\mathbb{Q}$ with the group of automorphic characters $\chi$ on $U\cong\mathbb{G}_{m}^{d}$ such that the induced character $(\chi_{\mathbf{m}})_{\mathbf{m}\in\Gamma_{M^{\circ}}}$ on $\mathbb{G}_{m}^{\Gamma_{M^{\circ}}}$ is identically $1$. For a place $v$, let $\mathcal{K}_{v}$ be the maximal compact subgroup of $\mathbb{G}_{m}(K_{v})^{d}$: i.e. for each prime number $p$ we have $\mathcal{K}_{p}=\mathbb{G}_{m}(\mathbb{Z}_{p})^{d}$ and $\mathcal{K}_{\infty}=\{\pm 1\}^{d}\subset\mathbb{G}_{m}(\mathbb{R})^{d}$. For any place $v\in\Omega_{\mathbb{Q}}$ given element $x\in(\mathcal{X},\mathcal{M})(\mathcal{O}_{v})$ and any element $t\in\mathcal{K}_{v}$, $tx$ also lies in $(\mathcal{X},\mathcal{M})(\mathcal{K}_{v})$ and
@@
which implies that $\hat{\tau}(\chi)$ is zero unless $\chi$ is identically $1$ on where $\mathcal{K}=\prod_{v\in\Omega_{\mathbb{Q}}}\mathcal{K}_{v}$. Therefore, in order to determine the Tamagawa constant, we only need to consider characters in the group $B$, where $B\subset\operatorname{Br}_{1}(X,M)/\operatorname{Br}\mathbb{Q}$ is the subgroup of characters which are trivial on $\mathcal{K}$. By class field theory [25, Ch. VI, §6; Ch. VII, §6], the group of automorphic characters on $\mathbb{G}_{m}$ which are both torsion and trivial on the maximal compact subgroup is isomorphic to the class group of $\mathbb{Q}$. Thus as the class group of $\mathbb{Q}$ is trivial, $B$ only contains the trivial character, proving that only the trivial character contributes to the leading constant in [24, Conjecture 1.4]. ∎
@@
[14, Théorème 1] Let $f\colon\mathbb{N}^{k}\rightarrow\mathbb{R}$ be a nonnegative arithmetic function and let $F$ be the associated Dirichlet series
@@
Theorem 3.11 has a typo in its original formulation [14, Théorème 1], as it requires the the upper bound in (P3) to be satisfied on the whole of $\mathscr{D}(\delta_{1}-\epsilon^{\prime})$. This is a stronger assumption than necessary and desired, as it would imply that $H$ is bounded on $\mathbb{R}_{>0}^{k}$, which would exclude the original application of the theorem in [13].
@@ 3.5.2. Choice of $\boldsymbol{\alpha}$
Recall that a linear form $N\rightarrow\mathbb{Z}$ is irreducible if it is not a positive multiple of another linear form. For $\tau$ a facet of $\sigma$, we write $u_{\sigma,\tau}\in N^{\vee}$ for the unique irreducible linear form which is zero on $\tau$ and positive on $\sigma\setminus\tau$. This linear form is simply given as $u_{\sigma,\tau}=n^{*}_{\rho_{i}}$, where $\rho_{i}$ is the unique ray in $\sigma(1)\setminus\tau(1)$. We also write $a_{\sigma,\tau}=a_{i}$. By [30, Lemma 8.9(i)] there is a unique maximal cone $\sigma^{\prime}\neq\sigma$ containing $\tau$, which satisfies $\tau=\sigma^{\prime}\cap\sigma$. Since $u_{\sigma^{\prime},\tau}$ is also irreducible and zero on $\tau$, we see $u_{\sigma^{\prime},\tau}=\pm n_{\rho_{i}}^{*}$. Since there are $n_{1}\in\sigma\setminus\tau$, $n_{2}\in\sigma^{\prime}\setminus\tau$ such that $n_{1}+n_{2}\in\tau$, we see $u_{\sigma^{\prime},\tau}=-n_{\rho_{i}}^{*}=-u_{\sigma,\tau}$.
@@ 3.5.3. Choice of linear forms
Now, as in de la Bretèche’s work [13, §4.3], we will use the upper bound
@@
valid for all $\epsilon>0$ and $T>0$, which follows from [33, Theorem II.3.8]. By shrinking $\delta_{1}$ if necessary, we can assume that $G$ extends to a holomorphic function on the topological closure of $\boldsymbol{\alpha}+\mathscr{D}(\delta_{1})$. Now since $G(\mathbf{s})$ is bounded on $\boldsymbol{\alpha}+\mathscr{D}(\delta_{1})$, this implies that condition (P3) in Theorem 3.11 is satisfied with $\delta_{2}=1/3$.
@@ 3.5.4. Determining the rank
where $A$ is the matrix such that the coefficient in position $(i,j)$ is $\langle\mu_{D^{\prime}}(\sigma_{j}),n_{\rho_{i}}\rangle$. Since a nef divisor on a toric variety is globally generated by [12, Theorem 6.3.12], we see by [12, Theorem 6.1.7] that $\operatorname{rank}(A)=\dim P_{L}$, where $P_{L}$ is the polyhedron associated to $L$ as defined in [12, §4.3]. By [12, Lemma 9.3.9] we also have $\dim P_{L}=\dim(X)$, as $L$ is big. The vector space $V$ is contained in the kernel of $\operatorname{Div}_{T}(X)_{\mathbb{Q}}\rightarrow\operatorname{Pic}(X)_{\mathbb{Q}}$. By [12, Theorem 4.2.1.] this kernel has dimension $\dim(X)$, but $\operatorname{rank}(l^{(1)},\dots,l^{(n)})=\dim(X)+1$ implies $V$ has dimension $\dim(X)$ as well, so $V$ is equal to the kernel. ∎
@@
The constant $C_{\infty}$ in Theorem 3.3 does not depend on the choice of the integers $d_{1},\dots,d_{n}$ determining the pair $(X,\overline{M^{\circ}})$ as in Notation 2.31. Therefore, we will assume that the integers are chosen large enough to ensure $a(\mathbb{Q},(X,M),L)=a(\mathbb{Q},(X,\overline{M^{\circ}}),L)$. Note that such integers exist as $(X,M^{\circ})$ is quasi-proper with respect to $L$.
@@
[14, Théorème 2(ii), Remarques (ii)] In the setting of Theorem 3.11, assume that the following additional conditions are satisfied:
@@
Now it remains to estimate the volume of the set $D(B)$. In order to simplify notation, we assume $M=M^{\circ}$, which we can do without loss of generality as the definition of $D(B)$ only depends on $(X,M^{\circ})$ and not on $(X,M)$. The set $D(B)$ is a generalization of the set $D(B)$ defined by Salberger [30, Notation 11.28] in his study of rational points on split toric varieties, and we will use the same approach he used to estimate its volume.
We regard $D(B)$ as a closed subset of the real locus of the universal torsor of $(X,M)$, where the universal torsor is as in Definition 2.34. As we assumed $M=M^{\circ}$ and $a((X,M),L)=1$, we have $\operatorname{pr}^{*}_{M}D=-D_{(X,M)}$. In order to estimate the volume of $D(B)$, Salberger splits it up as $D(B)=\cup_{\sigma\in\Sigma_{\max}}D(B,\sigma)$ using what he calls the toric canonical splitting [30, Notation 11.31]. We will similarly split up $D(B)$, but we will use the splitting induced by the fan $\Sigma_{\overline{M}}$ as given in Notation 2.31, rather than $\Sigma$. The fan $\Sigma_{\overline{M}}$ has the property that for all $\mathbf{m}\in\Gamma_{M}$, the ray spanned by an element $\phi(\mathbf{m})\in N$ lies in $\Sigma_{\overline{M}}$, and all rays in $\Sigma_{\overline{M}}$ are of this form. This property will aid in computing $D(B)$.
@@
As in [30, Proposition 11.22], we can describe when a point lies in $C_{M,\sigma,0}(\mathbb{R})$. The cone $\sigma$ contains exactly $d=\dim X$ rays. Let $d_{1}$ be the number of rays in $\sigma$ which lie in $\Sigma_{M}$. Let $r=\#\Sigma_{M(\sigma)}(1)-d$ be the number of raysin $\Sigma_{M(\sigma)}$ which lie outside of $\sigma$. Order the rays $\rho_{1},\dots,\rho_{r+d}$ in $\Sigma_{M(\sigma)}$ such that $\rho_{r+1},\dots,\rho_{r+d_{1}}$ are the rays in both $\sigma$ and $\Sigma_{M}$ and $\rho_{r+1},\dots,\rho_{r+d}$ are the rays in $\sigma$. Let $\mathbf{m}_{1},\dots,\mathbf{m}_{r+d}\in\Gamma_{M(\sigma)}$ be the elements corresponding to the rays $\rho_{1},\dots\rho_{r+d}$, and set $n^{(i)}=\phi(\mathbf{m}_{r+i})$ for $i=1,\dots,d$. As $n^{(1)},\dots,n^{(d)}$ are integer multiples of the ray generators of $\rho_{r+1},\dots,\rho_{r+d}$, they freely generate a finite-index sublattice $N_{\sigma}$ of $N$. Let $(\mu^{(1)},\dots,\mu^{(d)})$ be the corresponding dual $\mathbb{Z}$-basis of $N_{\sigma}^{\vee}\supset N^{\vee}$ and set
@@
The proof is identical to the proof of [30, Proposition 11.22]. ∎
@@
Consequently, we find the following analogue of [30, Equation (11.37)]
@@
obtained by evaluating at the anticanonical class $K_{(X,M)}$ of $(X,M)$. As the volume of the fibre above any $y\in[0,\infty)$ is equal to $\operatorname{Pic}(X,M)_{\mathrm{torsion}}\alpha_{\mathrm{Peyre}}((X,M),L)y^{b(\mathbb{Q},(X,M),L)-1}$, where $\alpha_{\mathrm{Peyre}}((X,M),L)$ is the variant of the $\alpha$-constant as given in [24, Remark 8.6], integrating along the fibres of $\lambda$ gives
@@
Now the formula [3, Example 2.1] for the exponential integral over a cone gives